# Beyond Pixels: Enhancing LIME with Hierarchical Features and Segmentation Foundation Models

### v4

@@ 4.2 Hierarchical segmentation
The segmentation capabilities of foundation models like SAM, influenced by its design and hyperparameters, allow fine and coarse segmentation of an image (Kirillov et al., 2023). These models have the ability to segment a human-recognized concept at various levels, from the entirety of a car to its components, such as doors or windshields. This multitude of segments enables the composition of a concept into its sub-concepts, creating a hierarchical segmentation. We enhance the LIME framework by introducing hierarchical segmentation, allowing users to specify the granularity of the segment for more personalized explanations. The architecture enables the surrogate model to learn about features driven by human-recognizable concepts iteratively. DSEG starts by calculating the importance scores of the coarse segments in the first stage. The segments identified as highly important are subsequently refined into their finer components, followed by another importance score calculation. Next, we detail the steps involved in DSEG (as illustrated in Figure 2) to explain an image within the LIME framework, and Figure 3 shows the outputs of its intermediate steps. Additionally, the pseudocode for the proposed framework is presented in Section A.1 for clarity and reference.
Figure 3: Visualized DSEG pipeline. Image (1) serves as the initial input, leading to its automatic segmentation depicted in (2). The hierarchical tree generated from this segmentation is illustrated in (3b), and (3a) showcasing the mask composed of first-order nodes. Image (4) displays the finalized mask created after eliminating empty spaces, which is fed back into the sample generation of LIME. Image (5) represents the resultant explanation within the DSEG-LIME framework. The image shows an instance from the COCO dataset (Lin et al., 2014), classified as an ’Airliner’ ($p$ = 0.86) by EfficientNetB4. Node 23 (blue node) indicates the segment that represents the superpixel of the airliner.
@@
where $\mathbf{x}$ denotes the input image, and $G_{\text{prompt}}$ specifies a general prompt configuration designed to enable automated segmentation. The output, $M_{\text{auto}}$, represents the automatically generated mask, as shown in Figure 3 (2). For this work, we used SAM with a grid overlay, parameterized by the number of points per side, to facilitate the automated segmentation process.
@@
In this study, we set $\theta=500$ to reduce the feature set. The remaining superpixels in $\mathcal{S^{\prime}}$ are considered for feature attribution. We incorporate this feature into DSEG to enable user-driven segment exclusion during post-processing, giving users control over the granularity within the segmentation hierarchy. This ensures that users can tailor the segmentation to their specific needs, thereby enhancing the method’s flexibility and adaptability.
@@
The hierarchy prioritizes parent segments (e.g., person) over child segments (e.g., clothing), as depicted (3a) in Figure 3. Each node represents one superpixel with its unique identifier. Section A.2 presents the BuildHierarchy method in greater detail. The depth $d$ of the hierarchy determines the granularity of the explanation, as defined by the user. A new set $\mathcal{S}^{\prime}_{d}$, with $d=1$, includes all nodes below the root. For $d>1$, DSEG does not start from the beginning. Instead, it uses the segmentation hierarchy and segments $\mathcal{S^{\prime}}$ from the first iteration. It then adds the nodes of the children of the top $k$ (a user-defined hyperparameter) most significant parent nodes in $\mathcal{S}^{\prime}_{d}$ at depth $d-1$, identified during the feature attribution phase. We visualize this selection in the tree shown in Figure 3 (3b), where all nodes with depth one, including the children of node 23, are considered in the second iteration. For the scope of this paper, we concentrate on the first-order hierarchy ($d=1$) but provide additional explanations with $d=2$ in the Section B.8.
@@
Although this modifies the distinctiveness of concepts, it enhances DSEG-LIME’s explanatory power. DSEG then utilizes the features $s_{0},\ldots,s_{d}\in\mathcal{S}_{d}$ for feature attribution within LIME. Figure 3 (4) shows the corresponding mask along with the explanation of $d=1$ in step (5) for the ’airliner’ class. An ablation study of these steps is in Section C.1 and in Section C.6 we show exemplary feature attribution maps.
@@ 5 Evaluation
In the following section, we will outline our experimental setup (Section 5.1) and introduce the XAI evaluation framework designed to assess DSEG-LIME both quantitatively (Section 5.2) and qualitatively (Section 5.3), compared to other LIME methodologies utilizing various segmentation algorithms. Subsequently, we discuss the limitations of DSEG (Section 5.4).
@@ 5.1 Experimental setup
Segmentation algorithms. Our experiment encompasses, along with SAM (vit_h), four conventional segmentation techniques: Simple Linear Iterative Clustering (SLIC) (Achanta et al., 2012), Quickshift (QS) (Hoyer et al., 2019), Felzenszwalb (FS) (Felzenszwalb & Huttenlocher, 2004) and Watershed (WS) (Neubert & Protzel, 2014). We carefully calibrate the hyperparameters of these techniques to produce segment counts similar to those generated by SAM. This calibration ensures that no technique is unfairly advantaged due to a specific segment count – for instance, scenarios where fewer but larger segments might yield better explanations than many smaller ones. In the Appendix, we demonstrate the universal property of integrating other segmentation foundation models within DSEG by presenting additional experiments with DETR (Carion et al., 2020) and SAM 2 (Ravi et al., 2024) in Section B.6, Section B.7.
@@
Models to explain. The models investigated in this paper rely on pre-trained models, as our primary emphasis is on explainability. We chose EfficientNetB4 and EfficientNetB3 (Tan & Le, 2019) as the ones treated in this paper, where we explain EfficientNetB4 and use EfficientNetB3 for a contrastivity check (Nauta et al., 2023) (Section 5.2.1). To verify that our approach works on arbitrary pre-trained models, we also evaluated it using ResNet-101 (He et al., 2015; maintainers & contributors, 2016) (Section B.1), VisionTransformer (ViT-384) (Dosovitskiy et al., 2020) (Section B.2) and ConvNext (Tiny-224) (Liu et al., 2022)(Section B.3). Furthermore, we demonstrate the applicability of our approach on a zero-shot learning example of CLIP (Radford et al., 2021) using a new dataset with other classes (Section C.4).
@@
Dataset. We use images from the ImageNet classes (Deng et al., 2009), on which the covered models were trained (Tan & Le, 2019; He et al., 2015; Dosovitskiy et al., 2020). Our final dataset consists of 100 carefully selected instances (Section D.1), specifically chosen to comprehensively evaluate the techniques quantitatively. We aligned with this number to maintain consistency with the sample size used in previous studies (Tan et al., 2024; Sun et al., 2023). However, we want to emphasize that the selection of images is not biased toward any model. We also test the approach for another dataset in Section C.4.
Hyperparameters and hardware setup. The experiments were conducted on an Nvidia RTX A6000 GPU. We compare standard LIME, SLIME (Zhou et al., 2021), GLIME (Tan et al., 2024), and BayLIME (Zhao et al., 2020), all integrated with DSEG, using 256 samples per instance, a batch size of ten and mean superpixel value for perturbation. For each explanation, up to three features are selected based on their significance, identified by values that exceed the average by more than 1.5 times the standard deviation. In BayLIME, we use the ‘non-info-prior’ setting to avoid introducing priors benefiting one method, even though the original papers show that this can have a diminishing effect. For SAM, we configure it to use 32 points per side, and conventional segmentation techniques are adjusted to achieve a similar segment count, as previously mentioned. In SLIC, we modify the number of segments and compactness; in Quickshift, the kernel size and maximum distance; in Felzenszwalb, the scale of the minimum size parameter; and in Watershed, the number of markers and compactness. Other hyperparameters remain at default settings to ensure a balanced evaluation across methods.
@@
Table 1: Quantitative summary - classes. The table presents four quantitative areas and their metrics, comparing five segmentation techniques applied to EfficientNetB4: DSEG with SAM and comparative methods SLIC, Quickshift (QS), Felzenszwalb’s (FS), and Watershed (WS). We test each with four LIME framework variations: LIME (L), SLIME (S), GLIME (G), and BayLIME (B). The experimental setup and metrics are detailed in Section 5.1 and Section 5.2.1. The table shows class metrics, each with a max score of 100. Arrows indicate performance improvement, highlighting the best scores for each metric bold.
@@ 5.2 Quantitative evaluation
We adapt the framework by Nauta et al. (2023) to quantitatively assess XAI outcomes in this study, covering three domains: content, presentation, and user experience. In the content domain, we evaluate correctness, output completeness, consistency, and contrastivity. Presentation domain metrics like compactness and confidence are assessed under content for simplicity. We will briefly describe each metric individually to interpret the results correctly. The user domain, detailed in Section 5.3, includes a user study that compares our approach with other segmentation techniques in LIME. We use quantitative and qualitative assessments to avoid over-emphasizing technical precision or intuitive clarity (Molnar et al., 2022).
@@ 5.2.1 Quantitative metrics definition
Correctness involves two randomization checks. The model randomization check (Random Model) (Adebayo et al., 2020) tests whether changing random model parameters alters explanations, while the explanation randomization check (Random Expl.) (Luo et al., 2020) examines if random output variations in the predictive model produce different explanations. In Table 1, we count instances where reintroducing explanations leads to different predictions. The domain employs two further deletion techniques: single deletion (Albini et al., 2020) and incremental deletion (Hoyer et al., 2019; Goyal et al., 2019). Single deletion acts as an alternative metric for assessing explanation completeness by replacing less relevant superpixels with a background and observing their effect on model predictions (Ramamurthy et al., 2020). We then record instances where the model retains the correct classification. Incremental deletion (Incr. Deletion) involves progressively removing features from most to least significant based on explanatory importance. We track the model’s output changes and quantify the impact by measuring the area under the curve (AUC) of the model’s confidence as explanation parts are excluded. This process continues until a classification change is detected (not the ground truth class), with the mean AUC score reported in Table 2.
@@
Output completeness measures whether an explanation covers the crucial area for accurate classification. It includes a preservation check (Preservation) (Goyal et al., 2019) to assess whether the explanation alone upholds the original decision, and a deletion check (Deletion) (Zhang et al., 2023) to evaluate the effect of excluding the explanation on the prediction outcome (Ramamurthy et al., 2020). This approach assesses both the completeness of the explanation and its impact on the classification. The results are checked to ensure that the consistency of the classification is maintained. Compactness is also considered, highlighting that the explanation should be concise and cover all the areas necessary for prediction (Chang et al., 2019), reported by the mean value.
@@
Consistency assesses explanation robustness to minor input alterations, like Gaussian noise addition, by comparing pre-and post-perturbation explanations for stability against slight changes (Noise Stability) (Zhang et al., 2021; Bhatt et al., 2021), using both preservation and deletion checks. For consistency of the feature importance score, we generate explanations for the same instance 16 times (Rep. Stability), calculate the standard deviation $\sigma_{i}$ for each coefficient $i$, and then average all $\sigma_{i}$ values. This yields $\bar{\sigma}$, the average standard deviation of coefficients, and is reported as the mean score. Furthermore, we reference (Bora et al., 2024) and compute a Gini coefficient to assess the inequality of the coefficients of the superpixels. We also evaluate DSEG directly within SLICE in Section B.5.
@@
Contrastivity integrates several previously discussed metrics, aiming for target-discriminative explanations. This means that an explanation $e_{\textbf{x}}$ for an instance x from a primary model $f_{1}$ (EfficientNetB4) should allow a secondary model $f_{2}$ (EfficientNetB3) to mimic the output of $f_{1}$ as $f_{1}(\textbf{x})\approx f_{2}(e_{\textbf{x}})$ (Schwab & Karlen, 2019). The approach checks the explanation’s utility and transferability across models, using EfficientNetB3 for preservation and deletion tests to assess consistency.
5.2.2 Quantitative evaluation results
Table 1 presents the outcomes of all metrics associated with class-discriminative outputs. The numbers in bold signify the top results, with an optimal score of 20. We compare LIME (L) (Ribeiro et al., 2016) with the LIME techniques discussed in Section 2, SLIME (S) (Zhou et al., 2021), GLIME (G) (Tan et al., 2024), and BayLIME (B) (Zhao et al., 2020) in combination with DSEG and the segmentation techniques from Section 5.1. The randomization checks in the correctness category confirm that the segmentation algorithm bias does not inherently affect any model. This is supported by the observation that most methods correctly misclassify when noise is introduced or the model’s weights or predictions are shuffled. In contrast, DSEG excels in other metrics, surpassing alternative methods regardless of the LIME technique applied. In the output completeness domain, DSEG’s explanations more effectively capture the critical areas necessary for the model to accurately classify an instance, whether by isolating or excluding the explanation. This efficacy is supported by the single deletion metric, akin to the preservation check but with a perturbed background. Moreover, noise does not compromise the consistency of DSEG’s explanations. The contrastivity metric demonstrates DSEG’s effectiveness in creating explanations that allow another AI model to produce similar outputs in over half of the cases and outperform alternative segmentation approaches. Overall, the impact of the LIME feature attribution calculation remains relatively consistent, as we possess an average of 21.07 segments in the dataset under evaluation.
@@ 5.2.2 Quantitative evaluation results
Table 2: Quantitative summary - numbers. The table summarizes metrics from Section 5.2.1, focusing on those quantified by rational numbers like incremental deletion, compactness, representational stability, and average computation time across the examples, detailed in Section 5.1. Once more, the arrows denote the direction of improvement performance.
@@
Table 2 further illustrates DSEG’s effectiveness in identifying key regions for model output, especially in scenarios of incremental deletion where GLIME outperforms. Bolded values represent the best performance in their according category. Although compactness metrics show nearly uniform segment sizes across the techniques, smaller segments do not translate to better performance in other areas. Repeated experimentation suggests that stability is less influenced by the LIME variant (with the here chosen hyperparameters) and more by the segmentation approach. In addition, the Gini value highlights that the feature values are more distinct compared to others. Further experiments show that DSEG outperforms the other techniques regarding stability as the number of features increases. This advantage arises from the tendency of data-driven approaches to represent known objects uniformly as a single superpixel (Section C.3). Thus, if a superpixel accurately reflects the instance that the model in question predicts, it can be accurately and effortlessly matched with one or a few superpixels - such accurate matching leads to a more precise and more reliable explanation. Conventional segmentation algorithms often divide the same area into multiple superpixels, creating unclear boundaries and confusing differentiation between objects. The segmentation phase is the main differentiator regarding computation time; DSEG has longer processing times than the others (except Quickshift).

### v5

@@ 4.2 Hierarchical segmentation
The segmentation capabilities of foundation models like SAM, influenced by its design and hyperparameters, allow fine and coarse segmentation of an image [28]. These models can segment a human-recognized concept at various levels, from the entirety of a car to its components, such as doors or windshields. This multitude of segments enables the composition of a concept into its sub-concepts, creating a hierarchical segmentation.
Our proposed framework, DSEG, operates in successive stages. To harness this hierarchy within LIME, we introduce a depth parameter $d$ that allows users to specify the desired level of segmentation granularity for more personalized explanations. At $d=1$, we compute the importance scores for all segments (coarse) of the top level. Segments with top $k$ scores, a user‐defined value, are then subdivided into their immediate children, and at $d=2$ we recompute importance scores on these finer segments. This refine–and–rescore cycle continues, incrementing $d$ at each iteration, until the user’s maximum depth is reached or no further subdivisions exist. By allowing the surrogate model to "zoom in" iteratively on the most salient human‐recognizable parts, DSEG produces explanations that are both semantically rich and tailored to the user’s needs. Next, we detail each step of the DSEG pipeline (2), illustrate its intermediate outputs (3), and present the complete pseudocode for clarity (Section˜A.1).
Figure 3: Visualized DSEG pipeline. Image (1) serves as the initial input, leading to its automatic segmentation depicted in (2). The hierarchical tree generated from this segmentation is illustrated in (3b), and (3a) showcases the mask composed of first-order nodes. Image (4) displays the finalized mask created after eliminating empty spaces, which is fed back into the sample generation of LIME. Image (5) represents the resultant explanation within the DSEG-LIME framework. The image shows an instance classified as an ’Airliner’ ($p$ = 0.86) by EfficientNetB4. Node 23 (blue node) indicates the segment that represents the superpixel of the airliner.
@@
where $\mathbf{x}$ denotes the input image, and $G_{\text{prompt}}$ specifies a general prompt configuration designed to enable automated segmentation. The output, $M_{\text{auto}}$, represents the automatically generated mask, as shown in Figure˜3 (2). For this work, we used SAM with a grid overlay, parameterized by the number of points per side, to facilitate the automated segmentation process.
@@
In this study, we set $\theta=500$ to reduce the feature set (see Section˜C.1). The remaining superpixels in $\mathcal{S^{\prime}}$ are considered for feature attribution. We incorporate this feature into DSEG to enable user-driven segment exclusion during post-processing, giving users control over the granularity within the segmentation hierarchy. This ensures that users can tailor the segmentation to their specific needs, thereby enhancing the method’s flexibility and adaptability.
@@
The hierarchy prioritizes parent segments (e.g., person) over child segments (e.g., clothing), as depicted (3a) in Figure˜3. Each node represents one superpixel with its unique identifier. Section˜A.2 presents the BuildHierarchy method in greater detail. The depth $d$ of the hierarchy determines the granularity of the explanation, as defined by the user. A new set $\mathcal{S}^{\prime}_{d}$, with $d=1$, includes all nodes below the root. For $d>1$, DSEG does not start from the beginning. Instead, it uses the segmentation hierarchy and segments $\mathcal{S^{\prime}}$ from the first iteration. It adds the child nodes of the top-$k$ most significant parent nodes in $\mathcal{S}’_{d}$ at depth d1d-1, identified during feature attribution. These are the only nodes used to train the surrogate model. We visualize this selection in the tree shown in Figure˜3 (3b), where all nodes with depth one, including the children of node 23, are considered in the second iteration. For the scope of this paper, we concentrate on the first-order hierarchy ($d=1$) but provide additional explanations with $d=2$ in Section˜B.8.
@@
Although this modifies the distinctiveness of concepts, it enhances DSEG-LIME’s explanatory power. DSEG then utilizes the features $s_{0},\ldots,s_{d}\in\mathcal{S}_{d}$ for feature attribution within LIME. Figure˜3 (4) shows the corresponding mask along with the explanation of $d=1$ in step (5) for the ’airliner’ class. An ablation study of these steps is in Section˜C.1 and in Section˜C.7 we show exemplary feature attribution maps.
@@ 5 Evaluation
In the following section, we will outline our experimental setup (Section˜5.1) and introduce the evaluation framework designed to assess DSEG-LIME both quantitatively (Section˜5.2) and qualitatively (Section˜5.3), compared to other LIME methodologies utilizing various segmentation algorithms. Subsequently, we discuss the limitations of DSEG (Section˜5.4)
@@ 5.1 Experimental setup
Segmentation algorithms. Our experiment encompasses, along with SAM (vit_h), four conventional segmentation techniques: Simple Linear Iterative Clustering (SLIC) [1], Quickshift (QS) [23], Felzenszwalb (FS) [16] and Watershed (WS) [41]. We carefully calibrate the hyperparameters (e.g., kernel size, marker placement, number of segments, minimum segment size) to produce segment counts comparable to those generated by SAM. This calibration ensures that no technique is unfairly advantaged due to a specific segment count, for instance, scenarios where fewer but larger segments might yield better explanations than many smaller ones. In the supplementary material, we demonstrate the universal property of integrating other segmentation foundation models within DSEG by presenting additional experiments with DETR [8] and SAM 2 [48] in Section˜B.6, Section˜B.7.
@@
Models to explain. The models investigated in this paper rely on pre-trained models, as our primary emphasis is on explainability. We chose EfficientNetB4 and EfficientNetB3 [53] as the ones treated in this paper, where we explain EfficientNetB4 and use EfficientNetB3 for a contrastivity check [40] (Section˜5.2.1). To verify that our approach works on arbitrary pre-trained models, we also evaluated it using ResNet-101 [22, 36] (Section˜B.1), VisionTransformer (ViT-384) [14] (Section˜B.2), and ConvNext (Tiny-224) (Section˜B.3)[33]. Furthermore, we demonstrate the applicability of our approach on a zero-shot learning example of CLIP [44] using a new dataset with other classes (Section˜C.5).
@@
Dataset. We use images from the ImageNet classes [13], on which the covered models were trained [14, 22, 53]. Our final dataset consists of 100 carefully selected instances (Section˜D.1), specifically chosen to comprehensively evaluate the techniques quantitatively. We aligned with this number to maintain consistency with the sample size used in previous studies [52, 54]. However, we want to emphasize that the selection of images is not biased toward any model. We also test the approach for another dataset in Section˜C.5.
@@
Table 1: Quantitative summary - classes. The table presents four quantitative areas and their metrics, comparing five segmentation techniques applied to EfficientNetB4: DSEG with SAM and comparative methods SLIC, Quickshift (QS), Felzenszwalb’s (FS), and Watershed (WS). We test each with four LIME framework variations: LIME (L), SLIME (S), GLIME (G), and BayLIME (B). The experimental setup and metrics are detailed in Section˜5.1 and Section˜5.2.1. The table shows class metrics, each with a max score of 100. Arrows indicate performance improvement, highlighting the best scores for each metric bold.
@@
Hyperparameters and hardware setup. The experiments were conducted on an Nvidia RTX A6000 GPU. We compare standard LIME, SLIME [61], GLIME [54], and BayLIME [60], all integrated with DSEG, using 256 samples per instance, a batch size of ten and mean superpixel value for perturbation. For each explanation, up to three features are selected based on their significance, identified by values that exceed the average by more than 1.5 times the standard deviation. In BayLIME, we use the ‘non-info-prior’ setting to avoid introducing priors benefiting one method, even though the original papers show that this can have a diminishing effect. For SAM, we configure it to use 32 points per side, and conventional segmentation techniques are adjusted to achieve a similar segment count, as previously mentioned. In SLIC, we modify the number of segments and compactness; in Quickshift, the kernel size and maximum distance; in Felzenszwalb, the scale of the minimum size parameter; and in Watershed, the number of markers and compactness. Other hyperparameters remain at default settings to ensure a balanced evaluation across methods. Additional comparisons with SLICE [7] and EAC [52] are in the supplementary material.
@@ 5.2 Quantitative evaluation
We adapt the framework by Nauta et al. [40] to quantitatively assess XAI outcomes in this study, covering three domains: content, presentation, and user experience. In the content domain, we evaluate correctness, output completeness, consistency, and contrastivity. Presentation domain metrics like compactness and confidence are assessed under content for simplicity. We will briefly describe each metric individually to interpret the results correctly. The user domain, detailed in Section˜5.3, includes a user study that compares our approach with other segmentation techniques in LIME. We use quantitative and qualitative assessments to avoid over-emphasizing technical precision or intuitive clarity [38].
@@ 5.2.1 Quantitative metrics definition
Correctness involves two randomization checks. The model randomization parameter check (Random Model) [2] tests if changing the random model parameters leads to different explanations. The explanation randomization check (Random Expl.) [35] examines if random output variations in the predictive model yield various explanations. For both metrics, in Table˜1 we count the instances where explanations result in different predictions when reintroduced into the model under analysis. The domain also utilizes two deletion techniques: single deletion [3] and incremental deletion [21, 23]. Single deletion serves as an alternative metric to assess the completeness of the explanation, replacing less relevant superpixels with a specific background to evaluate their impact on the model predictions [45]. After these adjustments, we note instances where the model maintains the correct image classification. Incremental deletion (Incr. Deletion) entails progressively eliminating features from most to least significant based on their explanatory importance. We observe the model’s output variations, quantifying the impact by measuring the area under the curve (AUC) of the model’s confidence, as parts of the explanation are excluded. This continues until a classification change is observed (not ground truth class), and the mean AUC score for this metric is documented in Table˜2.
@@
Output completeness measures whether an explanation covers the crucial area for accurate classification. It includes a preservation check (Preservation) [21] to assess whether the explanation alone upholds the original decision, and a deletion check (Deletion) [59] to evaluate the effect of excluding the explanation on the prediction outcome [45]. This approach assesses both the completeness of the explanation and its impact on the classification. The results are checked to ensure that the consistency of the classification is maintained. Compactness is also considered, highlighting that the explanation should be concise and cover all the areas necessary for prediction [9], reported by the mean value.
@@
Consistency assesses explanation robustness to minor input alterations, like Gaussian noise addition, by comparing pre-and post-perturbation explanations for stability against slight changes (Noise Stability) [58], using both preservation and deletion checks. For consistency of the feature importance score, we generate explanations for the same instance 16 times (Rep. Stability), calculate the standard deviation $\sigma_{i}$ for each coefficient $i$, and then average all $\sigma_{i}$ values. This yields $\bar{\sigma}$, the average standard deviation of coefficients, and is reported as the mean score. Furthermore, we reference [7] and compute a Gini coefficient to assess the inequality of the coefficients of the superpixels. We also evaluate DSEG directly within SLICE in Section˜B.5.
@@
Contrastivity integrates several previously discussed metrics, aiming for target-discriminative explanations. This means that an explanation $e_{\textbf{x}}$ for an instance x from a primary model $f_{1}$ (EfficientNetB4) should allow a secondary model $f_{2}$ (EfficientNetB3) to mimic the output of $f_{1}$ as $f_{1}(\textbf{x})\approx f_{2}(e_{\textbf{x}})$ [51]. The approach checks the explanation’s utility and transferability across models, using EfficientNetB3 for preservation and deletion tests to assess consistency.
@@
Table 2: Quantitative summary - numbers. The table summarizes metrics from Section˜5.2.1, focusing on those quantified by rational numbers like incremental deletion, compactness, representational stability, and average computation time across the examples, detailed in Section˜5.1. Once more, the arrows denote the direction of improvement performance.
@@
5.2.2 Quantitative evaluation results
Table˜1 presents the outcomes of all metrics associated with class-discriminative outputs. The numbers in bold signify the top results, with an optimal score of 100. We compare LIME (L) [49] with the LIME techniques discussed in Section˜2, SLIME (S) [61], GLIME (G) [54], and BayLIME (B) [60] in combination with DSEG and the segmentation techniques from Section˜5.1. The randomization checks in the correctness category confirm that the segmentation algorithm bias does not inherently affect any model. This is supported by the observation that most methods correctly misclassify when noise is introduced or the model’s weights or predictions are shuffled. In contrast, DSEG excels in other metrics, surpassing alternative methods regardless of the LIME technique applied. In the output completeness domain, DSEG’s explanations more effectively capture the critical areas necessary for the model to accurately classify an instance, whether by isolating or excluding the explanation. This efficacy is supported by the single deletion metric, akin to the preservation check but with a perturbed background. Moreover, noise does not compromise the consistency of DSEG’s explanations. The contrastivity metric demonstrates DSEG’s effectiveness in creating explanations that allow another AI model to produce similar outputs in over half of the cases and outperform alternative segmentation approaches. Overall, the impact of the LIME feature attribution calculation remains relatively consistent, as we possess an average of 21.07 segments in the dataset under evaluation.
Table˜2 further illustrates DSEG’s effectiveness in identifying key regions for model output, especially in scenarios of incremental deletion where GLIME outperforms. Bolded values represent the best performance in their according category. Although compactness metrics show nearly uniform segment sizes across the techniques, smaller segments do not translate to better performance in other areas. Repeated experimentation suggests that stability is less influenced by the LIME variant (with the chosen hyperparameters) and more by the segmentation approach. In addition, the Gini value highlights that the feature values are more distinct compared to others. Further experiments show that DSEG outperforms the other techniques regarding stability as the number of features increases. This advantage arises from the tendency of data-driven approaches to represent known objects uniformly as a single superpixel (Section˜C.4). Thus, if a superpixel accurately reflects the instance that the model in question predicts, it can be accurately and effortlessly matched with one or a few superpixels - such accurate matching leads to a more precise and more reliable explanation. Conventional segmentation algorithms often divide the same area into multiple superpixels, creating unclear boundaries and confusing differentiation between objects. DSEG has an average runtime of approximately 30 seconds across different settings. Although the standard LIME approach using SLIC operates at around 16 seconds, the additional computational steps in DSEG (e.g., SAM integration and hierarchical ordering) introduce only a modest overhead. Importantly, when compared to the widely used Quickshift option, which typically takes about 29 seconds, DSEG’s runtime is directly comparable. Thus, the extra computational cost is minimal relative to the substantial benefits in explanation quality.
The supplementary experiments (see Appendix˜B) exhibit similarly consistent quantitative gains. We also extend our analysis by benchmarking against a SHAP‑based explainer [52], integrating DSEG into the SLICE framework [7], and testing different segmentation models.